derive equation of rectangular componets of torque

Grade:11

1 Answers

vikas askiitian expert
13 years ago

T=r*F

r=xi+yj+zk

F=Fxi+Fyj+Fzk

T=(xi+yj+zk)(Fxi+Fyj+Fzk)

  T(torque)=(yFz-zfy)i + (zFx-xFz)j +(xFy-yFx)k